# Sub Specie Aeternitatis: Fourier Transforms from the Theory of Heat to Musical Signals

### v2

@@ Abstract.
in which we see the characters’ in their formative years
@@ 9. Dirac Comb
which will be referred to as the Dirac comb. Similarly to the delta, we should refrain from it calling a function, and our use in this context will be simply as an operator in expressions (as demonstrated later). Heuristically, it can be interpreted as an impulse train.
@@ 11. Discrete Time and Discrete Frequency
We now have the means to go from continuous to discrete time in a more rigorous way. By reasoning that the process of forming a sampled17 discrete sequence of evenly-spaced values from a continuous function $x(t)$ can be equated to applying a set of deltas at each point, we can then apply a Dirac comb as an operator to represent this process18,
@@
This is effectively the result of the series discussed in §8 put in a more generic form. It tells us precisely that the sampled spectrum produces a periodic waveform with period $F^{-1}$. We should note that these equations do not imply that the underlying continuous-time waveform $x(t)$ has a spectrum that is band-limited to an interval $F$ or that the spectrum $X(f)$ corresponds to a waveform that is time-limited to an interval $T$. If they are not, there will be an overlap, in frequency or in time, between the repeated copies of the spectrum or waveform. This effect is called aliasing19
@@ 12. Finite Time and Finite Frequency
We can now apply this to our problem of switching on/off a waveform20,
@@
which corresponds to the spectrum21
@@ 13. Discrete Transforms
for $n\in\mathbb{Z}$. The two transforms can be used transparently, e.g. forwards and inverse, like the continuous versions, to recover the original sequence, but a scaling factor $1/N$ needs to be applied once in either direction22.
@@
Si nous considérons en effet un morceau contenant plusieurs mesures (ce qui est le moins qu’on puisse demander) et qu’une note, la par exemple, figure une fois dans le morceau, l’analyse harmonique nous présentera la fréquence correspondante avec une certaine amplitude et une certaine phase, sans localiser le la dans le temps. Or, il est évident qu’au cours du morceau il est des instants où l’on n’entend pas le la. La représentation est néanmoins mathématiquement correcte, parce que les phases des notes voisines du la sont agencées de manière à détruire cette note par interférence lorsqu’on ne l’entend pas et à la renforcer, également par interférence, lorsqu’on l’entend; mais s’il y a dans cette conception une habileté qui honore l’analyse mathématique, il ne faut pas se dissimuler qu’il y a également une défiguration de la réalité: en effet, quand on n’entend pas le la, la raison véritable est que le la n’est pas émis23. [14, p.64]
@@
where the time domain is represented by the complex function $\Psi(t)$24. As can be seen we now have a function of both (radian) frequency $f$ and time $t$.
@@
and utilising a complex sinusoid representation25 instead of a real signal to facilitate the analysis. This produces a spectral component which is also subject to an envelope, this time in frequency, as its dual,

### v3

@@ Abstract.
in which we see the characters in their formative years
@@ 9. Dirac Comb
which will be referred to as the Dirac comb, denoted by 17. Similarly to the delta, we should refrain from it calling a function, and our use in this context will be simply as an operator in expressions (as demonstrated later). Heuristically, it can be interpreted as an impulse train.
@@ 11. Discrete Time and Discrete Frequency
We now have the means to go from continuous to discrete time in a more rigorous way. By reasoning that the process of forming a sampled18 discrete sequence of evenly-spaced values from a continuous function $x(t)$ can be equated to applying a set of deltas at each point, we can then apply a Dirac comb as an operator to represent this process19,
@@
This is effectively the result of the series discussed in §8 put in a more generic form. It tells us precisely that the sampled spectrum produces a periodic waveform with period $F^{-1}$. We should note that these equations do not imply that the underlying continuous-time waveform $x(t)$ has a spectrum that is band-limited to an interval $F$ or that the spectrum $X(f)$ corresponds to a waveform that is time-limited to an interval $T$. If they are not, there will be an overlap, in frequency or in time, between the repeated copies of the spectrum or waveform. This effect is called aliasing20
@@ 12. Finite Time and Finite Frequency
We can now apply this to our problem of switching on/off a waveform21,
@@
which corresponds to the spectrum22
@@ 13. Discrete Transforms
for $n\in\mathbb{Z}$. The two transforms can be used transparently, e.g. forwards and inverse, like the continuous versions, to recover the original sequence, but a scaling factor $1/N$ needs to be applied once in either direction23.
@@
Si nous considérons en effet un morceau contenant plusieurs mesures (ce qui est le moins qu’on puisse demander) et qu’une note, la par exemple, figure une fois dans le morceau, l’analyse harmonique nous présentera la fréquence correspondante avec une certaine amplitude et une certaine phase, sans localiser le la dans le temps. Or, il est évident qu’au cours du morceau il est des instants où l’on n’entend pas le la. La représentation est néanmoins mathématiquement correcte, parce que les phases des notes voisines du la sont agencées de manière à détruire cette note par interférence lorsqu’on ne l’entend pas et à la renforcer, également par interférence, lorsqu’on l’entend; mais s’il y a dans cette conception une habileté qui honore l’analyse mathématique, il ne faut pas se dissimuler qu’il y a également une défiguration de la réalité: en effet, quand on n’entend pas le la, la raison véritable est que le la n’est pas émis24. [14, p.64]
@@
where the time domain is represented by the complex function $\Psi(t)$25. As can be seen we now have a function of both (radian) frequency $f$ and time $t$.
@@
and utilising a complex sinusoid representation26 instead of a real signal to facilitate the analysis. This produces a spectral component which is also subject to an envelope, this time in frequency, as its dual,